20. Show that 20 shows that Siſi-1). either converges or does not converge.
Answer #1

- (20) DAlembent Ratio Test: a (tive real numbers. Then lim Mitil ito ai Lut <a:> be sequence of if 11 q let, then I ai convlim ito Aiti ai (+1)(1+1-1) lim i to slin å ina (it) (i+1-1) - ith ů (1-1) = lim ((+1) ito 311-1) = lim í (+ /;) - in a 31/1-:. l = ☆ <1 L a comenges. å M8 converges Scanned with CamScanner